ASSESSMENT: Construct the probability distribution of the situation below:

Two balls are drawn in succession without replacement from an urn containing 5 white balls and 6

black balls. Let B be the random variable representing the number of black balls.

Construct the probability distribution of the random variable B.



1
Expert's answer
2021-05-17T16:03:56-0400

The possible values that "B" can take are "0,1," and "2."

Each of these numbers corresponds to an event in the sample space "S=\{ww, wb, bw, bb\}" of equally likely outcomes for this experiment:

"B=0" to "\{ww\}," "B=1" to "\{wb, bw\}," "B=2" to "\{bb\}."

The probability of each of these events, hence of the corresponding value of "B," can be found simply by counting, to give


"B=0:"

"P(B)=\dfrac{5}{5+6}\cdot\dfrac{5-1}{5+6-1}=\dfrac{2}{11}"

"B=1:"

"P(B)=\dfrac{5}{5+6}\cdot\dfrac{6}{5+6-1}+\dfrac{6}{5+6}\cdot\dfrac{5}{5+6-1}=\dfrac{6}{11}"

"B=2:"

"P(B)=\dfrac{6}{5+6}\cdot\dfrac{6-1}{5+6-1}=\dfrac{3}{11}"



"\def\arraystretch{1.5}
 \begin{array}{c:c}
 B & 0 & 1 & 2 \\ \hline
\\
 P(B) &\dfrac{2}{11} & \dfrac{6}{11} & \dfrac{3}{11} \\
 
\end{array}"

This table is the probability distribution of "B."